OpenStudy (anonymous):

What is the equation of the line passing through (-3,6) and is parallel to the line y=4x-2 in standard form Choices: x-4y=-18 x+4y=18 4x-y=-18 4x+y=18

terenzreignz (terenzreignz):

What does it mean, for two lines to be parallel?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

My question wasn't a yes-or-no question...

OpenStudy (anonymous):

no matter how far they extend, they will never meet

terenzreignz (terenzreignz):

Yeah, actually, that is true... but what does it say about their slopes?

OpenStudy (anonymous):

they're equal?

terenzreignz (terenzreignz):

Yup :) so, what is the slope of y = 4x - 2 ?

OpenStudy (anonymous):

4 -- 1 or 4?

terenzreignz (terenzreignz):

Very good... So, our missing line must also have a slope of 4, to be parallel to y = 4x - 2, correct?

OpenStudy (anonymous):

right!

terenzreignz (terenzreignz):

We know the slope of our line, \(\large \color{blue}4\) and we know it passes through the point \(\large (\color{red}{-3 } \ , \ \color{green}6)\) When we know the slope \(\large \color{blue}m\) of a line and we know it passes through a point \(\large (\color{red}h \ , \ \color{green}k)\) Then the equation of the line is *always* given by this form: \[\huge y = \color{blue}mx - \color{blue}m\color{red}h+\color{green}k\]

terenzreignz (terenzreignz):

You know our slope and the point, just plug in :)

OpenStudy (anonymous):

so when you plug in would you get y=4(x)-4(-3)+6 which would simplify to y=4x+12+6 y=4x+18 ? is this right so far? so then since it needs to be in standard form (ax+by=c) you would rearrange it to say 4x+y=18 which would be the answer?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

I don't see how y = 4x + 18 leads to 4x + y = 18

OpenStudy (anonymous):

Oh, I'm probably wrong but I thought you could simply just rearrange the equation around to fit the required formula

terenzreignz (terenzreignz):

yes, in particular, y = 4x + 18 It's the same as 4x + 18 = y just switch it around...

terenzreignz (terenzreignz):

Now, try arranging that in standard form...

OpenStudy (anonymous):

standard form is ax+by=c correct?

terenzreignz (terenzreignz):

Yes...

OpenStudy (anonymous):

So the answer in standard form would be 4x+y=18 ?

terenzreignz (terenzreignz):

I see what the problem is :) You forget that when you transfer terms between equal-signs, you have to CHANGE THE SIGN. Positive becomes negative, and negative becomes positive :)

OpenStudy (anonymous):

Ahhh! So would the left side of the equation remain the same and the positive 18 become -18?

terenzreignz (terenzreignz):

\[\Large x \color{green}{+c}= y\] If I want to move the c over to the right side, essentially, I'm subtracting it from both sides... \[\Large x\color{blue}{+c}\color{red}{-c}=y\color{red}{-c}\] \[\Large x = y\color{red}{-c}\]

terenzreignz (terenzreignz):

Let's start with this... \[\Large 4x + 18 = y\]

terenzreignz (terenzreignz):

What we want is the x and y on the left side, and the constant, 18, on the right. How to bring the (positive) y to the left side?

OpenStudy (anonymous):

you would need to subtract it, so when it is moved to the left side it would become negative

terenzreignz (terenzreignz):

Yes... so... \[\Large 4x +18\color{red}{-y }=0\] Now we want that +18 on the right side... what to do?

OpenStudy (anonymous):

subtract it to get it off the left side and onto the right, making it -18 on the right side.

terenzreignz (terenzreignz):

Correct :) \[\Large 4x -y = \color{blue}{-18}\] voila :)

OpenStudy (anonymous):

Which would make the equation 4x-y=-18, which is in standard form, and the right answer!!!! :D THANK YOU SO SO SO SO SO SO SO SO MUCH!!!!!!!!

terenzreignz (terenzreignz):

The particular skill of transferring terms from one side to another and changing the sign, it's called transposition... master that, and you'll master lines in no time :) Great job. *Almost* perfect. Perfect it next time, promise me, I know you can do it ^_^

OpenStudy (anonymous):

Thank you! :)